According to the Federal Communications Commission, 70% of all U.S. households have cable television. Erica selects a random sample of 1,000 households in Dupage Country, what is the probability of obtaining a random sample of at least 734 households with cable from a sample size 1,000?
Answer #1

Let X be the random variable denoting the number of

households having cable television.

Thus, X ~ Bin(1000, 0.7)

Hence, the required probability = P(X \geq 734)

= \sum_{x = 734}^{1000} \binom{1000}{x} (0.7)^x (1-0.7)^{1000-x} = 0.0098. (Ans).